1) Current ratio. 2) Average days to sell inventory. (Use average inventory.) 4] Return on 5) Gross identify which of answer and to c, which company appears to be charging higher prices for its goods? txplain your answer and identify which of the ratio(s) from Part1 rs to have the higher level of financial y appears to be the more efficient at using its assets? which of the ratios) from Part 1you used to reachyour conclusion.
Answer #1
Part-1
Calculation of ratios:
Particulars Formula Kroger Co Whole foods market
1 Current ratio Current ratio= Current assets/Current liabilities

Kroger= 8911/11408=2.08
Whole foods= 1544/1252= 2.78
0.78 1.23
2 Average days to sell inventory Average collection period=(Average Inventory*365)/Revenue

Average inventory:
Kroger= (6933+6801)/2=6867
Whole Foods= (500+441)/2=471

Average collection:
Kroger= 6867*365/108665=23.06
Whole foods=471*365/15389=11.17
23.06 11.17
3 Debt to asset Formula:
Total debt/ Total assets

Kroger= 25114/30356=0.83
Whole foods=1972/5741=0.34
0.83 0.34
4 Reurn on investment Earnings*100/ Investment

Kroger=2649/5442=48%
Whole foods=878/3769=23%
48% 23%
5 Gross margin percentage Gross profit*100/Revenue

Kroger= 22953*100/108665=21.12%
Whole foods= 5416*100/15389=35.19%
21.12% 35.19%
6 Asset turnover Sales/ Total assets

Kroger= 108665/30356=3.58
Whole foods= 15389/5741=2.68
3.58 2.68
7 Return on sales Earnings from contiuing operations/Revenue

Kroger= 2649*100/108665=24.38%
Whole Foods= 878*100/15389=5.71%
24.38% 5.71%
8 Plant assets to long term debt Plant-:
Kroger=17912
Whole foods=3163

Ratio=:
Kroger=17912/13711=1.31
Whole foods=3163/65=48.66
1.31 48.66

Part-2

Written Analysis:

  1. Kroger co appears to be more profitable by using Return on Sales ratio of which Kroger co. has higher percentage return of 24.38% as compared to whole foods market inc. which has only 5.71%. Also, return on investment of Kroger co is also high that is 48%.
  2. Kroger co. has higher financial risk since, its debt to asset ratio is 0.83 as compared to whole foods market inc which has 0.34.
  3. Whole foods market inc is charging higher prices for its goods since its gross margin percentage is higher that is 35.19% as compared to Kroger which has 21.12%.
  4. Kroger co. is more efficient in using its assets since its Assets turnover ratio is 3.58 which is higher than Whole foods market which has 2.68.